OpenStudy (anonymous):

8x + 5y = –5 7x – 3y = 3 solve for y.

OpenStudy (anonymous):

x=(-5y-5)/8 7((-5y-5)/8)-3y=3

OpenStudy (anonymous):

Solving the second equation \[x = 3 + 3y/7\] then substitute into the 1st equation \[8/7*(3 + 3y) + 5y = -5\] then solve for y \[24/7*3y + 5y = -5 - 24*3/7\] so y = 1

OpenStudy (anonymous):

Whoops I think y = -1